Becton Labs, Inc., produces various chemical compounds for industrial use. One compound, called Fludex, is prepared using an elaborate distilling process. The company has developed standard costs for one unit of Fludex, as follows:

Standard Quantity
or Hours
Standard Price
or Rate
Standard Cost
Direct materials 2.40 ounces $ 18.00 per ounce $ 43.20
Direct labor 0.70 hours $ 14.00 per hour 9.80
Variable manufacturing overhead 0.70 hours $ 3.00 per hour 2.10
Total standard cost per unit $ 55.10

During November, the following activity was recorded related to the production of Fludex:

  1. Materials purchased, 12,000 ounces at a cost of $198,000.
  2. There was no beginning inventory of materials; however, at the end of the month, 3,200 ounces of material remained in ending inventory.

  3. The company employs 20 lab technicians to work on the production of Fludex. During November, they each worked an average of 160 hours at an average pay rate of $12.00 per hour.

  4. Variable manufacturing overhead is assigned to Fludex on the basis of direct labor-hours. Variable manufacturing overhead costs during November totaled $4,800.

  5. During November, the company produced 3,600 units of Fludex.

Required:

1. For direct materials:

a. Compute the price and quantity variances.

b. The materials were purchased from a new supplier who is anxious to enter into a long-term purchase contract. Would you recommend that the company sign the contract?

2. For direct labor:

a. Compute the rate and efficiency variances.

b. In the past, the 20 technicians employed in the production of Fludex consisted of 5 senior technicians and 15 assistants. During November, the company experimented with fewer senior technicians and more assistants in order to reduce labor costs. Would you recommend that the new labor mix be continued?

3. Compute the variable overhead rate and efficiency variances.
Answer #1
Becton Labs
Variances
MPV = ( SP - AP ) * AQ purchased
MPV = ( 18 - 16.5 ) * 12000 = $ 18000 ( favourable ) AP = 198000/12000
MQV = ( SQ - AQ ) SP
MQV = ( 3600 * 2.4 - ( 12000 - 3200 )) * 18 = $ 2880 ( unfavourable ) AQ = 12000 - 3200
The new supplier is providing the materials at a price lower than the standard price but its quality seems to be poor
therefore, MPV is favourable but MQV is unfavourable.. But overall Material cost variance is favourable ,, therefore, the
offer from supplier can be accepted.
LRV = ( SR - AR ) AH
LRV = ( 14 - 12 ) 20 *160 = $ 6400 ( favourable)
LEV = (Sh - AH ) SR
LEV = ( 3600 * 0.7 - 3200 ) 14 = $ 9520 ( unfavourable)
No, I would not recommend that this policy should be continued because employing less skilled
people have resulted in cost savings but on the other hand resulted in poor performance as a result
efficiency variance turned out to be unfavourable and made Overall Labour cost variance
unfavourable .
VOHRV = ( SR - AR ) AH
VOHRV = ( 3 - 1.5 ) * 3200 = $ 4800 ( favourable ) AP = 4800 / 3200
VOHEV = ( SH - AH ) SR
VOHEV = ( 2520 - 3200 ) 3 = $ 2040 ( unfavourable)
Answer #2

SP SQ 18.00 Per ounce 2.4 Ounces 8640 Ounces 3600 Units 12000 Ounces 8800 Ounces 16.50 Per ounce AQP AQ AP SR Data Direct Mat18000 F 2880 U YES Answer Mateial Price Variance (AQ*AP)-(AQ*SP) Matrial Quanity Varaince (AQ*SP)-(SQ*SP) would you recommand-18000 2880 Numerical Calculation Mateial Price Variance ( 12000 * 16.50 ) - ( 12000 * 18.00 ) = Matrial Quanity Varaince ( 8